# Can Large Language Models Address Open-Target Stance Detection?

### v4

@@ Appendix A Prompting Details
Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality. To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets. Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ).
@@
You will be provided with a tweet, and your task is to generate a target for this tweet. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be 4 words. The output should only be the target, no other words. Do not provide any explanation but you MUST give an output, do not leave any output blank.
@@
Prompt Justification:
  • Clarity of Task: The task description explicitly states what the LLM needs to do—generate a target for a given tweet in TG prompt and for SD, it provides a clear and concise definition of stance classification. This ensures that the LLM understands the primary objective without ambiguity.
  • Focus on Relevance: By instructing the model that the target should be the topic of the tweet, it emphasizes the need for relevance, ensuring that the generated target accurately represents the tweet’s content.
  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Stance Options Provided: By specifying the possible stances (Favor, Against, None), the prompt sets clear categories for the model to choose from. This simplifies the models decision process and ensures the outputs are standardized.
  • Output Format Specification: Clearly specifying that the output should only be the target, with no additional words or explanations, helps to standardize the outputs, making them easier to evaluate.
  • Mandating Output Generation: By instructing the model not to leave any output blank and always to provide an answer, the prompt ensures robustness in responses, reducing the chance of missing data.
@@
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following tweet, generate the target for this tweet, and determine its stance towards the generated target. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be 4 words. If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. If the stance is in favor of the generated target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. The answer only has to be one of these three words: FAVOR, AGAINST, or NONE. Do not provide any explanation but you MUST give an output, do not leave any output blank. The output format should be: “‘Target: <target>, Stance: <stance>“‘.
Prompt Justification:
  • Integrated Task Definition: The prompt combines the definitions of both target generation and stance classification, ensuring the LLM understands the combined objective of generating a target and determining the stance towards it.
  • Clear and Concise Instructions: By clearly outlining the steps—first generating a target and then determining the stance towards it—the prompt guides the model through a logical sequence, reducing ambiguity and improving task performance.
  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Target Generation Constraints: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Stance Classification Options: Providing explicit stance options (Favor, Against, None) simplifies the decision process for the model and ensures consistency in the outputs.
  • Mandatory Output Generation: By instructing the model to always provide an output and not leave any fields blank, the prompt ensures robustness and completeness in the responses, reducing the chance of missing data.
  • Output Format Specification: Specifying the output format as ’Target: <target>, Stance: <stance>’ standardizes the responses, making them easier to parse and evaluate.
  • Focus on Simplicity and Relevance: The prompt emphasizes that only the stance should be returned as a single word and no additional text, which maintains simplicity and relevance in the outputs.
@@ Appendix B Model Descriptions
Llama-3-8B Instruct
@@
Mistral 7B Instruct
@@ Appendix C BTSD Metric Justification
Table 2 contains the results of the stratified experiments with different measure stance detection F1 score when giving as input the text and a target, varying the quality of this latter. The numbers are significantly different.
@@ Appendix D Hardware Setting
All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 4 cores, 32 GB of RAM and one NVIDIA V100 GPU.
Appendix E VAST Dataset Processing
@@ Appendix G BTSD Fine-tuning
Following TSE, we consider 4 datasets combined i.e. SemEval (Mohammad et al., 2016b), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li and Caragea, 2021), containing 19 targets, splitting samples by (70%-train, 10%-dev, 20%-test), maintaining equal distribution from each of the targets. To assess the quality of the generated targets, we train a stance classifier model (referred as BTSD) as follows. Given a text $x$ and a target $t$, we first formulate the input as a sequence $s=[[CLS]t[SEP]x]$ where $[CLS]$ is a token that encodes the sentence and $[SEP]$ is used to separate the sentence $x$ and the target $t$. Then the representation of $[CLS]$ token is used to predict the stance toward target $t$. Note that $t$ is the golden target here.
@@ Appendix H SemSim Details
Specifically, we employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words. After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
Appendix I OTSD with EZSTANCE
We explore the OTSD task using EZSTANCE (Zhao and Caragea, 2024), the largest zero-shot stance detection (ZSSD) dataset, to address how proprietary and open LLMs compare in both target generation (TG) and stance detection (SD) tasks when the target is explicitly or implicitly mentioned in the text. Given the open-target nature of our task, EZSTANCE was selected for its diverse range of domains and unique target set, reflecting real-world scenarios.
I.1 Dataset Processing
I.2 Result Analysis
As shown in Table 5, LLMs face challenges in target generation for non-explicit samples compared to when the target is explicitly mentioned, based on BTSD and HE scores. When comparing proprietary and open models, consistent with our earlier findings (§5.1), no single LLM consistently outperforms the others. While GPT-4o tends to generate higher-quality targets in the TG+SD approach, Mistral surpasses other models in some TG&SD cases. Across both explicit and non-explicit samples, all models perform noticeably better using the TG&SD approach.
For stance detection, as shown in 5, similar to our findings with TSE and VAST in Section 5.2, LLMs perform better in SC when the target is explicitly mentioned. When comparing proprietary and open models, GPT-4o and Mistral both demonstrate competitive performance across approaches. Overall, LLMs consistently achieve better stance detection in the TG&SD approach.
TG+SD TG&SD
Model SS BTSD HE SC SS BTSD HE SC
Explicit
[…]
Non-explicit
GPT-3.5 0.76 38.78 0.377 36.69 0.81 41.89 0.402 38.97
[…]
Table 5: Comparison of TG and SD performance on EZSTANCE dataset (Zhao and Caragea, 2024) among the LLMs using various metrics: TG is evaluated with SS (SemSim), BTSD (%), and HE (Human Evaluation) scores, while SC (%) measures SD performance. Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
Appendix J Explanation with Samples in TG and SD
J.1 Target Generation
@@ I.2 Result Analysis
Appendix K Human Annotator Agreement
@@
Dataset Description: The file includes a Tweet column with various Twitter posts. The Gold Target column represents the true target, indicating the primary topic or issue the tweet addresses, which in turn informs the stance noted in the Gold Stance column. For each tweet, there are eight generated targets (in columns T1 through T8) produced by different models, whose names remain anonymous for unbiased annotation. Additionally, there are empty cells (e.g., in the T1 Score column) designated for storing human evaluation scores corresponding to each generated target.
Definition of a Relevant Target: A generated target in the context of the target generation task is considered relevant if it accurately captures or closely aligns with the main topic, issue, or entity identified by the golden target in a given tweet. It should reflect the key subject matter or concern that the golden target represents, maintaining coherence with the overall context and meaning of the tweet.
Task Description: Your task is to assess the quality of each generated target by evaluating its relevance to the golden target in the tweet. You will do this by answering the following question and assigning a score to each generated target. Scoring criteria and examples are provided below.
“How closely does the generated target relate to the golden target in the tweet?”
@@
Metrics TG+SD TG&SD
Ex N-Ex Ex N-Ex
BTSD-TSE 38.47 30.73 37.88 32.14
BTSD-VAST 41.10 41.54 41.34 40.38
SC-TSE 47.36 38.11 49.14 36.00
SC-VAST 48.47 46.46 49.00 45.41
Table 7: Caption goes here TG (evaluated by BTSD) and SD (evaluated by SC) performance using COT prompting on both TSE and VAST datasets in both Explicit and Non-explicit cases across the two approaches: TG+SD, TG&SD. Ex and N-Ex represent Explicit and Non-explicit cases, respectively.
Appendix L COT Initial Result
We initiate an exploration of alternative prompting techniques by employing the chain-of-thought (COT) prompting strategy for both target generation (TG) and stance detection (SD) tasks, across two approaches: TG+SD and TG&SD. As an initial phase, in Table 7, we assess target quality using the BTSD metric and evaluate stance detection performance with the SC metric, applied to both the TSE and VAST datasets for explicit and non-explicit cases.
Appendix M Human Annotation Score Distribution
@@
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. The details on human annotator guideline are provided in Appendix K, and the annotator agreement is in Appendix Table 6.
Appendix N Vsiual Comparison of TG and SD Performance
To better illustrate the data in Table 1, we have plotted Figures 2 and 3, which represent the TG and SD performance, respectively. These figures provide a clearer visualization of the conclusions discussed in Sections 5.1 and 5.2.
Figure 2: Comparison of TG performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using various metrics: SS (Semantic Similarity), BTSD (%), and HE (Human Evaluation). The graphs are generated based on the data in Table 1.
Figure 3: Comparison of SD performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using SC (stance classification) score. The graphs are generated based on the data in Table 1.

### v5

@@ Appendix A Prompting Details
Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality. To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets. Based on our empirical study, we set the maximum target length to 5 for GPT and Gemini models and 4 for Llama and Mistral models to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ).
@@
You will be provided with a tweet, and your task is to generate a target for this tweet. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be #n words. The output should only be the target, no other words. Do not provide any explanation but you MUST give an output, do not leave any output blank.
@@
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following tweet, generate the target for this tweet, and determine its stance towards the generated target. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be #n words. If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. If the stance is in favor of the generated target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. The answer only has to be one of these three words: FAVOR, AGAINST, or NONE. Do not provide any explanation but you MUST give an output, do not leave any output blank. The output format should be: “‘Target: <target>, Stance: <stance>“‘.
@@ Appendix B Model Descriptions
Gemini-1.5-Flash-8B
is a decoder-only model with 8 billion parameters, designed for efficiency and speed. Part of the Gemini series, it incorporates optimizations tailored for rapid inference without compromising performance on diverse tasks.
Gemini-1.5-Pro
is an advanced decoder-only model in the Gemini series, featuring enhanced architectural modifications and fine-tuned parameters to handle complex multi-domain tasks. It is instruction-tuned and optimized for professional-grade applications.
Llama-3-8B-Instruct
@@
Llama-3-70B-Instruct
is a decoder-only model in the LLaMA (Touvron et al., 2023) series with 70 billion parameters. Building on the architectural advancements of Llama-2 (Touvron et al., 2023), it has been instruction-tuned to excel across a wide range of natural language processing tasks.
Mistral-small
@@
Mistral-Large
is a decoder-only model in the Mistral family, incorporating advanced features like Grouped-Query Attention and Sliding Window Attention. With 13 billion parameters, it builds upon the success of Mistral-small (Jiang et al., 2023) and has been instruction-tuned for improved task generalization.
@@ Appendix C BTSD Metric Justification
Table 2 contains the results of the stratified experiments with different measure stance detection F1 score when giving as input the text and a target, varying the quality of this latter. All the numbers in Table 2 are significantly different with p-values < 0.05 in both cases.
@@ Appendix D Hardware Setting
All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 8 cores, 32 GB of RAM and two NVIDIA A100 GPUs.
Appendix E Dataset Pre-processing
VAST
@@ Appendix G BTSD Fine-tuning
Following TSE, we consider 4 datasets combined i.e. SemEval (Mohammad et al., 2016b), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li and Caragea, 2021), containing 19 targets, splitting samples by (70%-train, 10%-dev, 20%-test), maintaining equal distribution from each of the targets. To assess the quality of the generated targets, we train a stance classifier model (referred as BTSD) as follows. Given a text $x$ and a target $t$, we first formulate the input as a sequence $s=[[CLS]t[SEP]x]$ where $[CLS]$ is a token that encodes the sentence and $[SEP]$ is used to separate the sentence $x$ and the target $t$. Then the representation of $[CLS]$ token is used to predict the stance toward target $t$. Note that $t$ is the golden target here. During the test phase, initially we considered three random seeds, but finally reported the scores with a single seed, as the score variance was low.
@@ Appendix H SemSim Details
We employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words. After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
Appendix I Explanation with Samples in TG and SD
I.1 Target Generation
@@ I.2 Stance Detection
Appendix J Human Annotation Process
Three graduate students, specializing in NLP, participated voluntarily as annotators in the evaluation of generated targets. They were chosen for their expertise and familiarity with the stance detection task, enabling them to effectively assess the relevance of the generated targets compared to the golden ones. First, we provided them with an annotation guideline (detailed in §J.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples. Each annotator then independently scored the targets generated by different LLMs for 500 randomly selected samples (300 explicit and 200 non-explicit). To assess inter-annotator agreement, we calculated Krippendorff’s $\alpha$ (Krippendorff, 2011) and Fleiss’s $\kappa$ (Fleiss, 1971), with the results presented in Table 5. The final score for each generated target was determined based on the majority vote among the annotators. In cases of a tie, the median score was used.
J.1 Annotation Guideline
@@
Dataset Description: The file includes a Text column with various Twitter posts or news comments. The Gold Target column represents the true target, indicating the primary topic or issue the tweet addresses, which in turn informs the stance noted in the Gold Stance column. For each tweet, there are twenty-four generated targets (in columns T1 through T24) produced by different models, whose names remain anonymous for unbiased annotation. Additionally, there are empty cells (e.g., in the T1 Score column) designated for storing human evaluation scores corresponding to each generated target.
Definition of a Relevant Target: A generated target in the context of the target generation task is considered relevant if it accurately captures or closely aligns with the main topic, issue, or entity identified by the golden target in a given text. It should reflect the key subject matter or concern that the golden target represents, maintaining coherence with the overall context and meaning of the text.
Task Description: Your task is to assess the quality of each generated target by evaluating its relevance to the golden target in the text. You will do this by answering the following question and assigning a score to each generated target. Scoring criteria and examples are provided below.
“How closely does the generated target relate to the golden target in the text?”
@@
TG+SD TG&SD
Metrics TSE VAST EZSTANCE TSE VAST EZSTANCE
Explicit
BTSD (GPT-3.5) 38.47 (38.43) 41.10 (41.67) 47.60 (49.27) 37.88 (39.60) 41.34 (44.25) 48.98 (49.70)
BTSD (GPT-4o) 40.63 (41.55) 41.59 (42.69) 48.12 (49.70) 41.31 (41.92) 42.33 (44.25) 48.76 (50.69)
SC (GPT-3.5) 37.91 (42.68) 47.47 (47.21) 40.12 (39.10) 44.86 (47.61) 47.00 (48.48) 40.74 (40.63)
SC (GPT-4o) 47.36 (44.78) 41.17 (40.77) 43.79 (45.93) 49.14 (46.83) 47.84 (49.38) 45.68 (46.22)
Non-explicit
BTSD (GPT-3.5) 30.73 (33.10) 41.54 (38.10) 40.76 (38.78) 32.14 (31.32) 40.38 (38.55) 40.54 (41.89)
BTSD (GPT-4o) 32.72 (35.14) 41.71 (38.55) 41.34 (42.78) 33.44 (36.12) 41.92 (39.84) 43.90 (45.70)
SC (GPT-3.5) 31.88 (32.66) 43.41 (45.54) 37.76 (36.69) 35.46 (33.94) 44.46 (45.80) 38.10 (38.97)
SC (GPT-4o) 36.00 (36.39) 38.27 (39.92) 38.34 (38.50) 38.11 (37.5) 42.81 (43.84) 37.54 (39.26)
Table 6: Preliminary results using the CoT prompting strategy with GPT-3.5 and GPT-4o models on three datasets (TSE, VAST, EZSTANCE), evaluated for both Target Generation (TG) using BTSD and Stance Detection (SD) using SC. Results in the parenthesis are from Table 1 following the "Task Definition" strategy. The results are presented for both Explicit (Ex) and Non-explicit (N-Ex) cases across two approaches: TG+SD and TG&SD.
Appendix K Chain-of-Thought (CoT) Results
We conduct a preliminary experiment using the CoT prompting strategy across both approaches: TG+SD and TG&SD in Table 6. Evaluating target generation and stance detection, we test the performance of CoT with GPT-3.5 and GPT-4o models and compare it to the scores achieved using the "Task Definition" strategy, as reported in Table 1. The results show that, based on the BTSD score, the "Task Definition" strategy generally outperforms CoT in target generation. However, for stance detection, evaluated using SC scores, the two strategies alternate in outperforming each other depending on the specific settings. The CoT prompt used in this experiment is as follows:
Prompt (TG):
Step 1: You will be provided with a text, think step by step, and explain what is the target of the text. A target should be the topic on which the text is talking.
Step 2: Therefore, based on your explanation, what is the final target of the text? The target can be a single word or a phrase, but its maximum length MUST be 5 words. The output should only be the target, NO OTHER WORDS or EXPLANATION.
Prompt (SD):
Step 1: Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following text, think step by step, and explain the stance (favor, against, or none) of the text towards the provided target.
Step 2: Therefore, based on your explanation, what is the final stance? If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. Only return the stance as a single word, NO OTHER WORDS or EXPLANATION.
Prompt (TG, SD):
Step 1: Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following text, think step by step, and explain what is the target of the text. A target should be the topic on which the text is talking. Based on the target explanation, think step by step, and explain the stance (favor, against, or none) of the text toward the target.
Step 2: Therefore, based on your explanation of the target and the stance, generate the target of the given text and determine the stance toward the generated target. The target can be a single word or a phrase, but its maximum length MUST be 5 words. If the stance is in favor of the target, write FAVOR; if it is against the target, write AGAINST; if it is ambiguous, write NONE. Only return the stance as a single word, NO OTHER WORDS or EXPLANATION. The final output format MUST be: “‘Target: <target>, Stance: <stance>’ ”.
Appendix L Human Annotation Score Distribution
@@
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. The details on human annotator guideline are provided in Appendix J.1, and the annotator agreement is in Appendix Table 5.